Determine the final temperature when 100 g of water at 20.0 °C mixes with 400 g of water at 80 °C and with 500 g of water at 60.0 °C 


1
Expert's answer
2021-02-28T06:15:09-0500

M1= 100g

T1=20°C

M2= 400g

T2= 80°C

M3= 500g

T3= 60°C

T= ?

Heat gained = Heat lost

M1(T-T1) = M2(T2-T) + M3(T3-T)

100(T-20) = 400(80-T) + 500(60-T)

100T + 400T + 500T = 32000 + 30000 + 2000

1000T = 64000

T= 64°C